[Ncep.nhc.nco_contacts] TSB Morning Rounds - Sat Sep 16, 2017
nhc.tsbadmin at noaa.gov
Sat Sep 16 13:35:54 UTC 2017
Here is today's summary of NHC computer operations:
--- Craig Mattocks
1. GOES-16 imagery stopped ingesting at 1415 UTC Friday due to a problem
with Harris Corporation's WxConnect (Java) software that halted the GOES-R
data stream. All sites (AWC, NHC, SPC) stopped receiving products via LDM
at 1430 UTC. Tech Control restarted all Harris services at these sites and
the data began flowing again at 1952 UTC.
1. ATCF issues: The problems with ATCF communicating with WCOSS on Friday
night has been resolved. No problems with the following 03 UTC cycle.
The network "flapping" (intermittently dropping connections between NHC's
ATCF servers and 2 of the 4 Tide login nodes) was caused by IBM
implementing a software patch on WCOSS during this period. NCO Linux
Sysadmin Doug Gaer diagnosed the cause and verified that the problem was
resolved at 21Z Friday.
2. ATCF issues: Hurricane Specialist Lixion Avila reported that the
performance of the ATCF is getting sluggish for Hurricane Jose. The ATCF
A-deck is currently 35 MB in size (approaching the final size of Irma's
A-deck, 39 MB).
Acting HSU Branch Chief Mike Brennan does not want to split the A-deck at
this time. The Hurricane Specialists can minimize this sluggishness by not
displaying certain analyses in the ATCF GUI.
With a hurricane (Jose) approaching the east coast of the US, a depression
(TD 14), and a wave in the Atlantic basin, plus a hurricane (Norma) and a
depression (TD 15) in the Pacific basin, NHC Hurricane Specialists are
quite busy and focused this Saturday morning.
All IT systems functioning normally. Key Messages for Jose appearing on NHC
-------------- next part --------------
An HTML attachment was scrubbed...
More information about the Ncep.nhc.nco_contacts